# fMRI predictors based on language models of increasing complexity recover brain left lateralization

### v1

@@ Abstract
Over the past decade, studies of naturalistic language processing where participants are scanned while listening to continuous text have flourished. Using word embeddings at first, then large language models, researchers have created encoding models to analyze the brain signals. Presenting these models with the same text as the participants allows to identify brain areas where there is a significant correlation between the functional magnetic resonance imaging (fMRI) time series and the ones predicted by the models’ artificial neurons. One intriguing finding from these studies is that they have revealed highly symmetric bilateral activation patterns, somewhat at odds with the well-known left lateralization of language processing. Here, we report analyses of an fMRI dataset where we manipulate the complexity of large language models, testing 28 pretrained models from 8 different families, ranging from 124M to 14.2B parameters. First, we observe that the performance of models in predicting brain responses follows a scaling law, where the fit with brain activity increases linearly with the logarithm of the number of parameters of the model (and its performance on natural language processing tasks). Second, we show that a left-right asymmetry gradually appears as model size increases, and that the difference in left-right brain correlations also follows a scaling law. Whereas the smallest models show no asymmetry, larger models fit better and better left hemispheric activations than right hemispheric ones. This finding reconciles computational analyses of brain activity using large language models with the classic observation from aphasic patients showing left hemisphere dominance for language.
@@ 1 Introduction
Given this state of affairs, one can only be surprised by the symmetric patterns highlighted by studies that have relied on computational models of language to predict fMRI brain time-courses in participants listening to naturalistic texts (e.g. Huth et al.,, 2016; Heer et al.,, 2017; Toneva and Wehbe,, 2019; Caucheteux and King,, 2022; Schrimpf et al.,, 2021; Pasquiou et al.,, 2023). For example, Huth et al., (2016) constructed word embeddings using a latent-semantic approach based on co-occurrence counts (Landauer and Dumais,, 1997) and used them as predictors of brain activity while participants listened to stories. The maps revealed by this approach were strikingly symmetric. This finding was replicated in subsequent studies that have used predictors derived from more advanced language models based on LSTM (Jain and Huth,, 2018) or Transformers (e.g. Toneva and Wehbe,, 2019; Schrimpf et al.,, 2021; Pasquiou et al.,, 2023).
In this paper, we use neural language models of increasing size (from GPT-2 with 124 million parameters to Qwen1.5-14B with 14.2 billion parameters) to fit fMRI data obtained from participants who listened to a naturalistic text. We find that a clear left-right asymmetry emerges as the size and performance of these models increases.
@@ 2 Methods
We use the publicly available fMRI dataset Le Petit Prince111 https://openneuro.org/datasets/ds003643/versions/2.0.1 which provides recordings from English, French and Chinese participants who had listened to the audiobook of Le Petit Prince in their native language while being scanned using functional magnetic resonance (TR=2 s; voxel size=3.75 $\times$ 3.75 $\times$ 3.8 mm). Technical details on fMRI acquisition and preprocessing can be found in the publication accompanying the dataset (Li et al.,, 2022). Here, we use data from 48 English speakers for whom fMRI data spatially normalized in the Montreal Neurological Institute space are available in the repository. All of them were right-handed according to the Edinburgh handedness questionnaire. For each participant, fMRI acquisition was divided into 9 runs lasting each for about 10 min.
@@
To reduce the computational burden of the study, and because we are interested in making inferences about the general population, we compute a group average from all subjects. In order to further reduce the computational cost of the study, all functional data are first resampled to 4 mm isotropic voxels, close to the original acquisition resolution. Before averaging, we compute a symmetric brain mask common to all subjects using nilearn compute_multi_epi_mask function with the threshold parameter set at 50%. The resulting mask, henceforth named whole brain volume, comprises 25877 voxels (1656 cm3). For all subjects and each run independently, voxels’ time-series are high-pass filtered with a cut-off of 128 s, linearly detrended and standardized (zero mean and unit variance). We then compute the average subject by taking the mean of all these values per voxel and per run. Finally, we trim the first 20 s and the last 20 s of each run, as these were found to present deviation artifacts, and standardize the resulting time-series.
@@
In order to evaluate the signal-to-noise ratio in each voxel, independently of any language model, we estimate the reliability of each voxel across participants. To do so, we split the group of all 48 subjects into two equal subgroups, compute the average response for each group and predict the BOLD time series from each voxel of one group from the activity of all the voxels from the other group. The fitting procedure follows the same method described in more details below when fitting brain response with neural network activations, and is based on a linear mapping from one set of responses to another, evaluated on a held-out run through cross-validation, and regularized using ridge regression (L2). For each run, for each voxel, we thus compute the correlation between the true activity and the activity predicted from the other group of subjects, with a linear model trained on the other runs. This procedure is repeated 10 times, using different splits of the subjects. The final inter-group correlation for each voxel is the average over these 10 trials, and is plotted in Fig. 1. The resulting map is quite consistent to the ones obtained in previous inter-subjects correlations language studies (e.g. Lerner et al.,, 2011). The plots presented in the paper are based on the subset of the 25% most reliable voxels, representing a total of 6469 voxels (414 cm3) – 3286 in the left hemisphere and 3183 in the right hemisphere. These voxels are in hot-colored regions delineated with the dashed lines on Fig. 1. Corresponding plots computed over whole brain data are provided as supplementary figures in appendix.
@@
We selected a set of a priori regions of interest from previous works on syntactic and semantic composition: the Temporal Pole (TP), anterior Superior Temporal Sulcus (aSTS) and posterior Superior Temporal Sulcus (pSTS) from Pallier et al., (2011), the Inferior Frontal Gyrus pars opercularis (BA44), triangularis (BA44) and orbitalis (BA47) from Zaccarella et al., (2017), and the Angular Gyrus/Temporal Parietal Junction (AG/TPJ) from Price et al., (2015). Each region was defined as a sphere of 10 mm radius centered on the following coordinates in the Montral Neurological Institute’s MNI152 space : TP (-48, 15, -27), aSTS (-54, -12, -12), pSTS (-51, -39, 3), AG/TPJ (-52, -56, 22), BA44 (-50, 12, 16), BA45 (-52, 28, 10), BA47 (-44,34,-8).
@@ 2.2 Encoding models
For each model, we separately consider the activity provided by each layer, plus the embedding layer. Hence, for a 12-layer model, we consider 13 layers. Let us notate $X_{l,k}$ the activity of a layer $l$ generated by the text of run $k$, (after convolution with the haemodynamic response function, as described above), and $Y_{v,k}$ the BOLD time series of voxel $v$ of run $k$. $X_{l,k}$ is a $n_{\text{scans},k}\times n_{\text{neurons}}$ matrix, and $Y_{v,k}$ a vector of dimension $n_{\text{scans},k}$, where $n_{\text{scans},k}$ is the number of scans of run $k$, and $n_{\text{neurons}}$ the number of dimensions of layer $l$ (for instance, gpt2-large model has 1280 dimensions, and Mistral-7B-v0.1 has 4096 dimensions). The goal is to predict the functional brain activity $Y_{v,k}$ using the artificial neural activity $X_{l,k}$ as regressors, simply using a linear mapping between the two: $Y_{v,k}=X_{l,k}\beta$, where $\beta$ is a vector of dimension $n_{\text{neurons}}$. In order to avoid overfitting, the $\beta$ coefficients are determined using ridge regression with regularization strength controlled by the parameter $\alpha$. More precisely, given a run $k$ used as test, all the remaining 8 runs are used as training set to determine the $\beta$ coefficients. The hyperparameter $\alpha$ is chosen using nested cross-validation among a range of possible values (16 values log-spaced between $10^{2}$ and $10^{7}$): among the runs used for training, one is used for validation, and the remaining 7 runs for training the ridge regression with each value of $\alpha$. Correlation between the actual values $Y_{v,\text{val}}$ and the predicted ones $\widehat{Y_{v,\text{val}}}$ is used to decide which value of $\alpha$ is used during training: the best $\alpha$ is chosen as the one yielding the greatest correlation averaged over all voxels  [^2] We performed exploratory analyses showing that a single value for $\alpha$ for all the voxels in the brain was not detrimental to the predictions of the functional data, and produced a massive gain in speed.. This value is then used to compute the ridge regression on the training runs and predict the functional brain activity on the test run $k$. The brain correlation for a given voxel $v$ and a given layer $l$ is given by repeating this procedure for all 9 runs and taking the average correlation: $1/n_{\text{runs}}\sum_{k=1}^{n_{\text{runs}}}\operatorname*{corr}(Y_{v,k};\widehat{Y_{v,k}})$. In the end, for a given voxel, we take as brain correlation the best correlation when considering all the layers of the model.
@@ 2.3 Computer code
The whole study took about 5 five days of CPU time on a computer equipped with an Intel Xeon w5-3425 processor (12 cores), mainly dedicated to perform the ridge regression of all layers of all models, using the scikit-learn package.
@@ 3.1 Distribution of correlations
Figure 2: Performance of various models in predicting fMRI brain time-courses. (a) Density estimates of the distributions of $r$-scores obtained for all 28 large language models, the random baselines and GloVe. The densities are scaled to have the same maximum (b) Average $r$-score as a function of the logarithm of model’s size. Here and in the next figures, the shaded area indicates the 95% confidence interval of the slope, computed with bootstrap. (c) Same, split by models’ family.
@@ 3.2 Scaling law of fMRI encoding models
Focusing on the contextual large language models, panel (b) of Fig. 2 shows the relationship between the size of the language models and their prediction performance, that is, the average correlation across the 25% most reliable voxels. The linear correlation between the brain correlation and the logarithm of the number of parameters is $r=0.95,p=8.5e-15$. The corresponding plot for the whole volume, displayed in panel (b) of  Fig. B.1, also shows a linear relationship ($r=0.95,p=3.1e-14$). This finding replicates the scaling law first described by Antonello et al., (2024).
The panel (c) of  Fig. 2 details the results by the family of models. Within each family (except Llama 2), the scaling law roughly holds. Older models like the ones from the OPT family perform worse than more recent ones like Mistral or Qwen, although within the same family the scaling law holds well in general, as exemplified by the GPT-2, OPT or Qwen1.5 families. Interestingly, the Mamba family, which is based on a recurrent neural network architecture, contrary to all the other models which are based on the Transformer decoder, has similar performances as encoding models of brain functional data, although it lies at the bottom of the envelope of the results (see Fig. B.2). In addition, we performed an analysis of the fit as a function of layer depth, depicted in Fig. B.3. The results confirms earlier reports by Toneva and Wehbe, (2019) and Caucheteux and King, (2022) that the middle layers are the most predictive ones.
@@
Fig. 3 shows the brain correlations maps associated to the two most extreme models: GPT-2 with the smallest number of parameters and Qwen1.5-14B with the largest number of parameters. Interestingly, while the map generated by GPT-2 is quite symmetric, Qwen1.5-14B fits fMRI responses better in the left hemisphere than in the right (while overall performing better than GPT-2).
@@
Figure 3: Brain correlation maps associated to the smallest (a) and the largest model (b). These maps show the increase in $r$-score relative the model using the random embedding baseline 1024d.
@@
The number of parameters of a model is a rough proxy of its performance on language tasks. For example, OPT-13B is found to have lower performances on various natural language processing tasks than Mistral-7B which has almost half the number of parameters. Hence, we also looked at measures of models’ performances beyond their raw number of parameters. First we compute the perplexity of all the models on the Wikitext-2 test set (Merity et al.,, 2016). Second we look at the Hellaswag benchmark (Zellers et al.,, 2019) which aims to measure the ability of a model to understand language. This measure was found to be quite challenging for language models, and the performance of models scale on this benchmark well with the training budget, either in terms of number of parameters or number of tokens seen during training (see Touvron et al.,, 2023). Fig. 5 shows how brain correlation increases with better models, either as measured by lower perplexity on the Wikitext-2 test set or on the Hellaswag benchmark. Both measures exhibit strong correlation with brain fit ($r=-0.91,p=9.5e-12$ for perplexity and $r=0.95,p=5.2e-15$ for Hellaswag). But this is particularly relevant when looking at the ten largest models, with number of parameters above 3 billions: whereas the correlation between brain score and number of parameters is no longer significant in this case ($r=0.28,p=0.4$), the other measures of performance display a significant correlation ($r=-0.85,p=1.7e-3$ for perplexity and $r=0.81,p=4.1e-3$ for Hellaswag): see Fig. 5.
@@ 3.3 Left-right hemispheric asymmetry
Voxel-based analyses.
@@
The maps presented on Fig. 3 and Fig. 4 hint at left-right differences. Therefore, we plot the relationships between model size and brain scores split by hemisphere in panel (a) of Fig. 6. This graphics reveals that the larger the model, the stronger the left-right asymmetry. More precisely, while the smallest models have similar $r$-scores both in the left and right hemisphere, the largest models yield stronger $r$-scores in the left one. The interaction between model size and hemisphere can be assessed by the correlation between model size and the left-right difference in $r$-scores, displayed in panel (b) of  Fig. 6; its $p$-value is $2.8e^{-10}$. This graphics shows that the difference also follows a scaling law. See Fig. B.5 for the same analysis on the whole brain volume.
Analyses in Regions of Interest.
The relationships between model size and $r$-scores in seven regions of interest from the language network in the left hemisphere, and in their mirror-image regions in the right hemisphere, are plot on Fig. 7. Brain scores improves with model size in all regions, both in the left and in the right hemisphere. Except for pars opercularis (BA44) in the inferior frontal gyrus, and for the temporal pole (TP), a left-right asymmetry in favor of the left emerges when model size increases. The steepest effect occurs in the Angular gyrus/Temporo-parietal. This region, AG_TPJ, was already highlighted in Fig. 4 showing the slopes in individual voxels.
@@
Figure 7: Model size impact in various Regions of Interest. (a) Locations of ROIs. (b) Slope in $r$-score as a function of the logarithm of the number of parameters, averaged over all voxels of a given ROI. Error bars indicate 95% confidence interval, estimated by bootstrap. The annotation below each ROI label indicates an estimate of the level of statistical significance of the difference between the slopes of the voxels of a left ROI compared to its respective right ROI (ns: non significant, *:$p<0.05$, **:$p<0.01$, ***∗ ∗ ∗:$p<0.001$). (c) Brain correlations over the random embedding baseline (see B.6 for similar plots with raw brain correlations), as a function of the logarithm of the number of parameters. The annotation next to each ROI label indicates the level of statistical significance of the correlation between the difference in $r$-scores of the models on the left hemisphere minus the right hemisphere, as a function of the logarithm of the number of parameters, testing the interaction between hemisphere and model size. See Table A.2 for full statistics.

### v2

@@ Abstract
Over the past decade, studies of naturalistic language processing where participants are scanned while listening to continuous text have flourished. Using word embeddings at first, then large language models, researchers have created encoding models to analyze the brain signals. Presenting these models with the same text as the participants allows to identify brain areas where there is a significant correlation between the functional magnetic resonance imaging (fMRI) time series and the ones predicted by the models’ artificial neurons. One intriguing finding from these studies is that they have revealed highly symmetric bilateral activation patterns, somewhat at odds with the well-known left lateralization of language processing. Here, we report analyses of an fMRI dataset where we manipulate the complexity of large language models, testing 28 pretrained models from 8 different families, ranging from 124M to 14.2B parameters. First, we observe that the performance of models in predicting brain responses follows a scaling law, where the fit with brain activity increases linearly with the logarithm of the number of parameters of the model (and its performance on natural language processing tasks). Second, although this effect is present in both hemispheres, it is stronger in the left than in the right hemisphere. Specifically, the left-right difference in brain correlation follows a scaling law with the number of parameters. This finding reconciles computational analyses of brain activity using large language models with the classic observation from aphasic patients showing left hemisphere dominance for language.
@@ 1 Introduction
Given this state of affairs, one can only be surprised by the symmetric patterns highlighted by studies that have relied on computational models of language to predict fMRI brain time-courses in participants listening to naturalistic texts (e.g. Huth et al.,, 2016; Heer et al.,, 2017; Toneva and Wehbe,, 2019; Caucheteux et al.,, 2021; Schrimpf et al.,, 2021; Pasquiou et al.,, 2023). For example, Huth et al., (2016) constructed word embeddings using a latent-semantic approach based on co-occurrence counts (Landauer and Dumais,, 1997) and used them as predictors of brain activity while participants listened to stories. The maps revealed by this approach were strikingly symmetric. This finding was replicated in subsequent studies that have used predictors derived from more advanced language models based on LSTM (Jain and Huth,, 2018) or Transformers (e.g. Toneva and Wehbe,, 2019; Caucheteux et al.,, 2021; Schrimpf et al.,, 2021; Pasquiou et al.,, 2023) (but see Caucheteux and King,, 2022 who reported a significant left-right asymmetry). One potential interpretation is that brain scores are essentially driven by semantic representations (Kauf et al.,, 2024), supposedly represented in a very distributed fashion across both hemispheres (see e.g. the discussion in Huth et al.,, 2016).
In this paper, we use 28 large language models (LLMs) of increasing size (from GPT-2 with 124 million parameters to Qwen1.5-14B with 14.2 billion parameters; see Table A.1 for the full list) to fit fMRI data obtained from participants who listened to a naturalistic text (from Le Petit Prince dataset, Li et al.,, 2022). We find that a clear left-right asymmetry emerges as the size and performance of these models increases, and that this is not simply due to differences in signal-to-noise ratio between left and right hemispheres.
@@ 2 Methods
We use the publicly available fMRI dataset Le Petit Prince111 https://openneuro.org/datasets/ds003643/versions/2.0.5 which provides recordings from English, French and Chinese participants who had listened to the audiobook of Le Petit Prince in their native language while being scanned using functional magnetic resonance (TR=2 s; voxel size=3.75 $\times$ 3.75 $\times$ 3.8 mm). Technical details on fMRI acquisition and preprocessing can be found in the publication accompanying the dataset (Li et al.,, 2022). Here, we use fMRI data spatially normalized in the Montreal Neurological Institute space, from all 49 English speakers. All of them were right-handed according to the Edinburgh handedness questionnaire. For each participant, fMRI acquisition was divided into 9 runs lasting each for about 10 min.
@@
To reduce the computational burden of the study, and because we are interested in making inferences about the general population, we compute a group average from all subjects. In order to further reduce the computational cost of the study, all functional data are first resampled to 4 mm isotropic voxels, close to the original acquisition resolution. Before averaging, we compute a symmetric brain mask common to all subjects using nilearn compute_multi_epi_mask function with the threshold parameter set at 50%. The resulting mask, henceforth named whole brain volume, comprises 25870 voxels (1656 cm3). For all subjects and each run independently, voxels’ time-series are high-pass filtered with a cut-off of 128 s, linearly detrended and standardized (zero mean and unit variance). We then compute the average subject by taking the mean of all these values per voxel and per run. Finally, we trim the first 20 s and the last 20 s of each run, as these were found to present deviation artifacts, and standardize the resulting time-series.
@@
In order to evaluate the signal-to-noise ratio in each voxel, independently of any language model, we estimate the reliability of each voxel across participants. To do so, we split the group of all 49 subjects into two (almost) equal subgroups (24/25), compute the average response for each group and predict the BOLD time series from each voxel of one group from the activity of all the voxels from the other group. The fitting procedure follows the same method described in more details below when fitting brain response with neural network activations, and is based on a linear mapping from one set of responses to another, evaluated on a held-out run through cross-validation, and regularized using ridge regression (L2). For each run, for each voxel, we thus compute the correlation between the true activity and the activity predicted from the other group of subjects, with a linear model trained on the other runs. This procedure is repeated 10 times, using different splits of the subjects. The final inter-group correlation for each voxel is the average over these 10 trials, and is plotted in Fig. 1. The resulting map is quite consistent to the ones obtained in previous inter-subjects correlations language studies (e.g. Lerner et al.,, 2011). The plots presented in the main paper are based on the subset of the 25% most reliable voxels, representing a total of 6468 voxels (414 cm3) – 3297 in the left hemisphere and 3171 in the right hemisphere. These voxels are in hot-colored regions delineated with the dashed lines on Fig. 1. In Fig. B.1, inter-subjects correlations averaged over parcels defined by the Harvard-Oxford atlas are plotted. The graphics in panels (b) and (c) reveal a strong relationship between correlations in the left and right homologous regions, with a tendency toward stronger correlations in the left than in the homologous right regions.
@@
We selected a set of a priori regions of interest from previous works on syntactic and semantic composition: the Temporal Pole (TP), anterior Superior Temporal Sulcus (aSTS) and posterior Superior Temporal Sulcus (pSTS) from Pallier et al., (2011), the Inferior Frontal Gyrus pars opercularis (BA44), triangularis (BA44) and orbitalis (BA47) from Zaccarella et al., (2017), and the Angular Gyrus/Temporal Parietal Junction (AG/TPJ) from Price et al., (2015). Each region was defined as a sphere of 10 mm radius centered on the following coordinates in the Montreal Neurological Institute’s MNI152 space : TP (-48, 15, -27), aSTS (-54, -12, -12), pSTS (-51, -39, 3), AG/TPJ (-52, -56, 22), BA44 (-50, 12, 16), BA45 (-52, 28, 10), BA47 (-44,34,-8).
@@ 2.2 Encoding models
For each model, we separately consider the activity provided by each layer, plus the embedding layer. Hence, for a 12-layer model, we consider 13 layers. Let us notate $X_{l,k}$ the activity of a layer $l$ generated by the text of run $k$, (after convolution with the haemodynamic response function, as described above), and $Y_{v,k}$ the BOLD time series of voxel $v$ of run $k$. $X_{l,k}$ is a $n_{\text{scans},k}\times n_{\text{neurons}}$ matrix, and $Y_{v,k}$ a vector of dimension $n_{\text{scans},k}$, where $n_{\text{scans},k}$ is the number of scans of run $k$, and $n_{\text{neurons}}$ the number of dimensions of layer $l$ (for instance, gpt2-large model has 1280 dimensions, and Mistral-7B-v0.1 has 4096 dimensions). The goal is to predict the functional brain activity $Y_{v,k}$ using the artificial neural activity $X_{l,k}$ as regressors, simply using a linear mapping between the two: $\widehat{Y_{v,k}}=X_{l,k}\beta$, where $\beta$ is a vector of dimension $n_{\text{neurons}}$. In order to avoid overfitting, the $\beta$ coefficients are determined using ridge regression with regularization strength controlled by the parameter $\alpha$. More precisely, given a run $k$ used as test, all the remaining 8 runs are used as training set to determine the $\beta$ coefficients. The hyperparameter $\alpha$ is chosen using nested cross-validation among a range of possible values (16 values log-spaced between $10^{2}$ and $10^{7}$): among the runs used for training, one is used for validation, and the remaining 7 runs for training the ridge regression with each value of $\alpha$. Correlation between the actual values $Y_{v,\text{val}}$ and the predicted ones $\widehat{Y_{v,\text{val}}}$ is used to decide which value of $\alpha$ is used during training: the best $\alpha$ is chosen as the one yielding the greatest correlation averaged over all voxels  [^2] We performed exploratory analyses showing that a single value for $\alpha$ for all the voxels in the brain was not detrimental to the predictions of the functional data, and produced a massive gain in speed.. This value is then used to compute the ridge regression on the training runs and predict the functional brain activity on the test run $k$. The brain correlation for a given voxel $v$ and a given layer $l$ is given by repeating this procedure for all 9 runs and taking the average correlation: $1/n_{\text{runs}}\sum_{k=1}^{n_{\text{runs}}}\operatorname*{corr}(Y_{v,k};\widehat{Y_{v,k}})$. In the end, for a given voxel, we take as brain correlation the best correlation when considering all the layers of the model.
@@ 2.3 Computer code
The main English study on the average subject takes about five days of CPU time on a computer equipped with an Intel Xeon w5-3425 processor (12 cores), mainly dedicated to perform the ridge regression of all layers of all models, using the scikit-learn package.
@@ 3.1 Distribution of correlations
Figure 2: Performance of various models in predicting fMRI brain time-courses. (a) Density estimates of the distributions of $r$-scores obtained for all 28 large language models, the random baselines and GloVe. The densities are scaled to have the same maximum (b) Average $r$-score as a function of the number of parameters of the model, in log scale. Here and in the next figures, the shaded area indicates the 95% confidence interval of the slope, computed with bootstrap. (c) Same, split by models’ family.
@@ 3.2 Scaling law of fMRI encoding models
Focusing on the contextual large language models, Fig. 2b shows the relationship between the size of the language models and their prediction performance, that is, the average correlation across the 25% most reliable voxels. The linear correlation between the brain correlation and the logarithm of the number of parameters is $r=0.95,p=7.3e-15$. The corresponding plot for the whole volume, displayed in panel (b) of  Fig. B.2, also shows a linear relationship ($r=0.95,p=3.2e-14$). This finding replicates the scaling law first described by Antonello et al., (2024).
Fig. 2c details the results by the family of models. Older models like the ones from the OPT family perform worse than more recent ones like Mistral or Qwen, although within the same family the scaling law holds well in general, as exemplified by the GPT-2, OPT or Qwen1.5 families. Interestingly, the Mamba family, which is based on a recurrent neural network architecture contrary to all the other, Transformer-based, models, has similar brain scores, although they lie at the bottom of the envelope of the results (see Fig. B.3). In addition, we performed an analysis of the fit as a function of layer depth, depicted in Fig. B.4. The results confirm earlier reports by Toneva and Wehbe, (2019) and Caucheteux and King, (2022) that the middle layers are the most predictive ones.
@@
Fig. 3 shows the brain correlations maps associated with the two most extreme models: GPT-2 with the smallest number of parameters and Qwen1.5-14B with the largest number of parameters. Interestingly, while the map generated by GPT-2 is quite symmetric, Qwen1.5-14B fits fMRI responses better in the left hemisphere than in the right (while overall performing better than GPT-2).
@@
Figure 3: Brain correlation maps associated with the smallest (a) and the largest model (b). These maps show the increase in $r$-score relative to the model using the random embedding baseline 1024d.
@@
The number of parameters of a model is a rough proxy of its performance on language tasks. For example, OPT-13B is found to have lower performances on various natural language processing tasks than Mistral-7B which has almost half the number of parameters. Hence, we also look at measures of models’ performance beyond their raw number of parameters. First we compute the perplexity of all the models on the Wikitext-2 test set (Merity et al.,, 2016). Second we look at the Hellaswag benchmark (Zellers et al.,, 2019), which aims to measure the ability of a model to understand language. This measure was found to be quite challenging for language models, and the performance of models scale on this benchmark well with the training budget, either in terms of number of parameters or number of tokens seen during training (see Touvron et al.,, 2023). Fig. 5 shows how brain correlation increases with better models, either as measured by lower perplexity on the Wikitext-2 test set or on the Hellaswag benchmark. Both measures exhibit strong correlation with brain fit ($r=-0.91,p=1.1e-11$ for perplexity and $r=0.95,p=4.2e-15$ for Hellaswag). But this is particularly relevant when looking at the ten largest models, with number of parameters above 3 billions: whereas the correlation between brain score and number of parameters is no longer significant in this case ($r=0.30,p=0.41$), the other measures of performance display a significant correlation ($r=-0.86,p=1.5e-3$ for perplexity and $r=0.83,p=3.0e-3$ for Hellaswag): see Fig. 5.
@@ 3.3 Left-right hemispheric asymmetry
Voxel-based analysis.
@@
The maps presented on Fig. 3 and Fig. 4 hint at left-right differences. Therefore, we plot the relationships between model size and brain scores split by hemisphere in Fig. 6a. This graphic reveals that the larger the model, the stronger the left-right asymmetry. More precisely, while the smallest models have similar $r$-scores both in the left and right hemisphere, the largest models yield stronger $r$-scores in the left one. The interaction between model size and hemisphere can be assessed by the correlation between model size and the left-right difference in $r$-scores, displayed in Fig. 6b. This graph shows that the difference also follows a scaling law ($r=0.89$, $p=1.8e^{-10}$). See Fig. B.6 for a similar analysis on the whole brain volume. Finally, as we did for the brain scores (Fig. B.4), we examined the effect of layer depth on the asymmetry. This analysis, reported on Fig. B.7, shows that the layers with the stronger fit also have the strongest left-right asymmetry.
Regions of Interest analysis.
The relationships between model size and $r$-scores in seven regions of interest from the language network in the left hemisphere, and in their mirror-image regions in the right hemisphere, are shown on Fig. 7. Brain scores improve with model size in all regions, both in the left and in the right hemisphere. Except for pars opercularis (BA44) in the inferior frontal gyrus, and for the temporal pole (TP), a left-right asymmetry in favor of the left emerges when model size increases. The steepest effect occurs in the Angular gyrus/Temporo-parietal. This region, AG_TPJ, was already highlighted in Fig. 4 showing the slopes in individual voxels.
@@
Figure 7: Model size impact in various Regions of Interest. (a) Locations of ROIs. (b) Slope in $r$-score as a function of the logarithm of the number of parameters, averaged over all voxels of a given ROI. Error bars indicate 95% confidence interval, estimated by bootstrap. The annotation below each ROI label indicates an estimate of the level of statistical significance of the difference between the slopes of the voxels of a left ROI compared to its respective right ROI (ns: non significant, *:$p<0.05$, **:$p<0.01$, ***:$p<0.001$). (c) Brain correlations over the random embedding baseline (see B.9 for similar plots with raw brain correlations), as a function of the logarithm of the number of parameters. The annotation next to each ROI label indicates the level of statistical significance of the correlation between the difference in $r$-scores of the models on the left hemisphere minus the right hemisphere, as a function of the logarithm of the number of parameters, testing the interaction between hemisphere and model size. See Table A.2 for full statistics.
Relationship between signal-to-noise ratio, brain scores, and growth in asymmetry.
Do large models fit the left hemisphere better than the right simply because the signal-to-noise ratio, or ceiling of explainable signal, is higher in the left voxels than in the right ones? A proxy for the explainable signal is the model-free inter-subject correlation (ISC). Fig. B.1 shows that the ISC indeed tends to be stronger on the left. To determine if this could explain our findings, we conducted several analyses presented on Fig. B.8 and Fig. B.9. First, we computed “normalized $r$-scores” by dividing raw $r$-scores by the ISC in each voxel. Panels (a) and (b) of Fig. B.8 display the relationship between these normalized brain correlations and model size. The left-right asymmetry previously observed with raw $r$-scores on Fig. 6 still holds. Next, using the parcels from the Harvard-Oxford atlas, we plot the relationship between ISC, brain correlations and their asymmetry (Fig. B.8, bottom panels). Although ISC and brain correlation correlate (panel c), the left-right difference in slope of brain correlation as a function of model size does not covary with the interhemispheric difference in ISC (panel d). Finally, we computed the ISC in our Regions of Interest. Fig. B.9 suggests that while differences in ISC partly explain the average brain score difference, they do not explain the increase in difference between left and right hemispheres, across models.
Impact of model training.
In order to assess the impact of model training on the left-right asymmetry, we compared $r$-scores for randomly initialized, untrained models, and for trained models, from the GPT-2 and the Qwen1.5 families. Fig. B.10a shows that, with training, brain correlations get better on the left hemisphere compared to the right one. Moreover, while the left-right score difference increases with the number of parameters for trained models, this relationship breaks down and is flat in the case of the untrained models. Another way to assess the influence of model training is to fit the same LLM at various stages of training. The Pythia family (Biderman et al.,, 2023) provides many intermediate checkpoints during training, from the untrained model, randomly initialized, to the full model, trained on one epoch on the training set of 300B tokens from the Pile dataset. Here we consider Pythia-6.9B, with 32 layers and 4096 dimensions. Fig. B.10b shows the impact of training on brain correlations. The model’s fit improves as training proceeds and the left-right asymmetry, originality biased towards the right hemisphere, increases, reaching a plateau favoring the left hemisphere.
3.4 Extension to other languages
As left dominance is universal (see e.g. Malik-Moraleda et al.,, 2022; it is even attested in sign language, Poizner et al.,, 1987), we wanted to check if our result holds for other languages. We therefore applied our approach to Chinese and French data provided in Le Little Prince dataset (considering all 35 Chinese participants and all 28 French participants).
We first computed inter-subjects correlations. Their distributions and the corresponding maps are displayed on Fig.B.11 and Fig.B.12 panels (a) and (b). The correlation values are lower than for English (Fig.B.1), but the most reliable voxels encompass the same brain regions. Panels (d) and (e) reporting correlations in regions from the Harvard-Oxford atlas (shown in panel (c)) and their left-right difference qualitatively replicate the previous observations in English results, that is, left-right asymmetries are detected in the same regions.
Although much fewer LLMs are available than for English, some models on the Hugging Face hub were trained on Chinese or French data. We considered 10 models for Chinese, ranging from 12M to 14B parameters, and 4 models for French, ranging from 124M to 7B parameters (see panel (f)). The relationships between model size and $r$-scores are presented in panels (g), (h) and (i) of Fig. B.11 and Fig. B.12. They confirm the scaling law, and the emergence of the left-right hemispheric asymmetry in two languages different from English.
3.5 Individual analyses
The analyses presented above were performed on data averaged across individuals. It is important to verify if the emergence of asymmetry holds at the individual level or is an artifact of averaging. Because of limitations in time and computational power available to us, we could only analyze five participants on a subset of models. To avoid any bias, we selected the first 5 English participants. The results, presented in  Fig. B.13, show that three of them present a significant increase in asymmetry ($p<0.05$) when considering the full brain (panel c), and four when considering the 10% voxels with the highest $r$-scores in each hemisphere (panel e).